### v1

# EDEN: Efficient Dual-Layer Exploration Planning for Fast UAV Autonomous Exploration in Large 3-D Environments

Abstract
Efficient autonomous exploration in large-scale environments remains challenging due to the high planning computational cost and low-speed maneuvers. In this paper, we propose a fast and computation-efficient dual-layer exploration planning method. The insight of our dual-layer method is efficiently finding an acceptable long-term region routing and greedily exploring the target in the region of the first routing area with high speed. Specifically, the proposed method finds the long-term area routing through an approximate algorithm to ensure real-time planning in large-scale environments. Then, the viewpoint in the first routing region with the lowest curvature-penalized cost which can effectively reduce decelerations caused by sharp turn motions will be chosen as the next exploration target. To further speed up the exploration, we adopt an aggressive and safe exploration-oriented trajectory to enhance exploration continuity. The proposed method is compared to state-of-the-art methods in challenging simulation environments. The results show the proposed method outperforms other methods in terms of exploration efficiency, computational cost, and trajectory speed. We also conduct real-world experiments to validate the effectiveness of the proposed method. The code will be open-sourced [^1] https://github.com/NKU-MobFly-Robotics/EDEN.
Index Terms:
Autonomous exploration, motion planning, trajectory planning.
I Introduction
Umanned aerial vehicles (UAVs) are popular in many 3-D exploration relevant applications, e.g., search and rescue [1, 2], environment inspection [3, 4], environment reconstruction [5, 6, 7, 8], among others. However, most existing methods remain low efficiency, low trajectory speed, and high computational cost.
Existing exploration methods can be coarsely separated into two kinds of methods: greedy methods and optimal-routing-based methods. Greedy methods attempt to select the target that maximizes the reward in a short period. The original frontier-based method [9] selects the closest frontier, i.e., the boundary voxel of explored and unknown space as an exploration target. To improve the UAV speed, [10] propose to select a frontier in camera field-of-view (FOV) to minimize the velocity change. By combining Next-Best-Views and rapid random tree (RRT), receding-horizon-next-best-views (RH-NBVP) [11] is proposed, and its variant methods [12, 13, 14, 15] show high exploration efficiency in small environments. However, these greedy methods suffer from massive back-and-forth maneuvers in large-scale environments due to the lack of long-term consideration. Another type of method, optimal-routing-based methods [16, 17, 18, 19, 20, 21], make decisions by constructing and solving routing problems to minimize the total routing length. The work in [21] proposes to find a coverage path by solving the traveling salesman problem (TSP). The method proposed by [16] constructs and solves the asymmetrical traveling salesman problem (ATSP) to find the shortest routing of frontiers. To further reduce back-and-forth maneuvers, [18] increases the routing priority of isolated frontiers. By solving routing problems, these methods dramatically reduce back-and-forth maneuvers. However, the computational cost grows fast with the number of routing nodes, which makes these methods have to execute stops for planning in large-scale environments. What is more, the work in [22] shows that the speed of UAVs in exploration still has great potential to be stimulated. Therefore, a computation-efficient, far-sighted, and high-speed method is expected.
Motivated by the above drawbacks of existing methods, we propose an Efficient Dual-layer Exploration plaNning method (EDEN) which can plan efficient exploration motions in hundred-meter-level 3-D environments in real-time. Different from greedy methods and optimal-routing-based methods, the proposed dual-layer method first finds a long-term sub-optimal exploration region routing efficiently and then decides the short-term targets that minimizes the exploration motion cost greedily. Specifically, we improve the Christofides algorithm, which is an approximate algorithm for solving TSP, to an exploration-oriented heuristic Christofides (EOHC) algorithm to find a sub-optimal but satisfying tour for ATSP with low computational cost. Then, a viewpoint for safety and two viewpoints with the highest curvature-penalized exploration gain in the first exploration region of the routing will be selected for high-speed trajectory planning. Then, inspired by [23], we propose an aggressive and safe exploration-oriented (ASEO) trajectory for high-speed continuous exploration. The ASEO trajectory is composed of three parts: an exploring part, a continuous part, and a safety part. The time of the exploring part, the time of the continuous part, and the feasibility of the safety part are optimized together to find an aggresive trajectory and a backup trajectory for safety.
We evaluate the proposed method by comparing it against state-of-the-art methods in various challenging environments. The results show that the proposed method has $5.2\%-14.0\%$ higher exploration efficiency, $10.4\%-15.5\%$ faster UAV speed, and $28.9\%-74.5\%$ lower computational cost. Further, the proposed method is deployed on a UAV and is tested in real-world environments to evaluate its effectiveness. The contributions of this work are summarized as follows:
  1. 1.
    We propose an exploration-oriented heuristic Christofides (EOHC) algorithm for efficiently solving large-scale ATSP. The proposed EOHC can find a satisfactory region routing with a low computational cost. The real-time performance of EOHC can be guaranteed even in hundred-meter-level environments.
  2. 2.
    To take advantage of the current UAV speed and enhance exploration continuity, we propose a heuristic curvature-penalized continuous gain for the short-term exploration targets choosing.
  3. 3.
    To make the UAV keep high-speed continuous exploration in unknown environments, we propose an ASEO trajectory planning method that ensures trajectory efficiency and safety at the same time.
  4. 4.
    Massive comparative simulations are conducted to evaluate the proposed method. We also test the proposed method by real-world experiments. We will release our source code.
II related work
II-A Greedy Exploration Planning
The frontier-based method is first proposed by Yamauchi et al. [9]. It selects the frontier voxel that is closest to the robot as the exploration target. To minimize the UAV velocity change, [10] selects the frontier voxel that is located at the boundary of camera FOV as the exploration target. Different from the greedy frontier-based methods that only consider the next step exploration, Bircher et al. [11] propose RH-NBVP which samples a set of exploration motion sequences and executes the most informative sequence’s first step and it shows high efficiency in short-term exploration. Nevertheless, in large-scale environments, it suffers from sample failure. To cope with this problem, Selin et al. [24] use frontier to guide sampling when a local sample fails. However, long-distance sampling is time-consuming and it does not guarantee success even with the guidance of frontier. To reduce the planning cost and improve long-distance exploration, roadgraphs, which can compress the complex explorable space information and tarversable space information into compact nodes and edges, are widely used in recent proposed greedy methods [12, 25, 14, 26, 27, 28, 29, 30]. Consequently, roadgraphs enable greedy methods of efficient planning in large environments. Although these methods show high efficiency in short-term exploration and low computational cost, they suffer from back-and-forth maneuvers due to a lack of long-term consideration.
II-B Optimal-Routing-Based Exploration Planning
Optimal-routing-based methods usually formulate the exploration target decision problems into routing problems. Optimal-tour-based methods can be roughly devided into frontier-ATSP methods [16, 17, 31, 18] and coverage-path-ATSP methods [32, 20, 21, 5, 19]. Frontier-ATSP methods formulate ATSP using frontiers and find the shortest tour between frontiers. Coverage-path-ATSP methods formulate ATSP with unexplored space and find the shortest coverage path. The main drawback of the optimal-routing-based methods is that they have to spend luxury computation on solving ATSP in large-scale environments. Moreover, counterintuitively, recent works [28, 33] report that optimal-tour-based methods usually do not have an obvious advantage in minimizing the exploration path distance compared to greedy methods. This is because optimal-routing-based methods have implicit optimal assumptions, where frontier-ATSP methods assume a new frontier will not be generated after exploring an existing frontier, while coverage-path-ATSP methods assume all the unexplored space is traversable. Obviously, the above assumptions do not hold in most cases. Therefore, spending too much computation on finding a high-quality ATSP solution is not cost-effective.
II-C Exploration-Oriented Trajectory Planning
Exploration trajectory plays a crucial role in exploration tasks. The early proposed exploration methods [11, 9] stop-and-go motions are used. To utilize the good maneuverabilities of UAVs, Dharmadhikari et, al. [15] adopts motion primitives to speed up the exploration. To further improve the trajectory speed, Zhou et al. [16] optimizes minimum-time trajectory with optimizable terminal velocity and acceleration for high-speed exploration. Although minimum-time trajectory is very efficient, the optimizable terminal velocity and acceleration do not ensure being beneficial to continuously exploring the newly generated frontier after exploring the previous frontier. What is more, the optimizable terminal velocity and acceleration can lead to collision.
In summary, a far-sighted and computationally efficient exploration planning method and a trajectory planning method that enable fast continuous exploration while ensuring safety are required for exploration in large-scale environments.
Figure 1: The overview of the proposed method framework.
III system overview
The proposed method is composed of a dual-layer exploration target decision module that can plan both the long-term region routing and the short-term viewpoint target, and a trajectory planning module that can plan high-speed continuous exploration trajectories as shown in Fig. 1. A dynamic voxel map with resolution $r_{v}$ is updated by integrating depth image and odometry information. Different from the normal voxel map, the dynamic voxel map only maintains a local part of voxels around the UAV while other voxels are dynamically read and written to save memory usage in large environments. Based on the voxel map, a dynamic topological graph (DTG), is maintained to manage exploration-related information. The traversable information is maintained by DTG’s history nodes and edges between them, and the explorable space information is maintained by DTG’s explorable region of interest (EROI), which is a set of uniform scale cubic regions that can be explored by visiting corresponding active viewpoints. For detailed information, we refer readers to our previous work [34]. The exploration planning utilizes DTG to efficiently formulate ATSP, and ATSP is solved through EOHC to find satisfactory exploration region routing with small computation. Then, three viewpoints, including an exploring viewpoint, a continuous viewpoint, and a safety viewpoint in the first region of the region routing are found for the following trajectory planning. Then, an aggressive and safe exploration-oriented trajectory which separates at the exploring viewpoint, and its two branches end at a continuous viewpoint and a safe viewpoint correspondingly, is optimized. The exploring part of the aggressive and safe exploration-oriented trajectory will be executed.
IV Dual-layer exploration viewpoints planning
In this section, we first propose the EOHC algorithm for finding a satisfactory region routings efficiently in large-scale environments. Then, we use curvature-penalized continuous gain to decide the targets in the first region of the tour for the following high-speed ASEO trajectory planning.
Figure 2: An instance of EOHC-based routing planning, the numbers indicate the lengths of edges. (a) An MST that connects all the explorable history nodes of DTG is extracted. (b) The leaf point, whose distance is the longest to the root node, and its corresponding branch (green) will be executed last. (c) Once, there are multiple siblings for the current node, the sibling (orange dashed curve) whose first node has the shortest distance to the current node will be chosen as the next branch. (d) The final routing found by the proposed EOHC-based routing planning.
IV-A EOHC-Based Global Routing Planning
As analyzed in Section. II.B, the optimal routing of frontiers or unexplored space is very likely not the optimal routing for exploration due to the implicit optimal assumption not being met. Therefore, instead of finding a high-quality of all precise explorable frontiers, the proposed method aims to spend low computation on finding a satisfactory coarse routing of explorable regions. The process of EOHC-based routing planning is shown in Fig. 2.
Firstly, we extract all the history nodes of DTG that connect at least one EROI as the explorable regions. To reduce the computation when formulating ATSP, the distances between all explorable regions and the distances between the robot and all explorable regions are computed by conducting the Dijkstra algorithm on DTG. Thanks to the sparse DTG, the ATSP can be formulated efficiently.
To solve ATSP efficiently, we design an EOHC algorithm, which is improved from the Christofides algorithm [35] and tailored to ATSP. The key idea of the Christofides algorithm is to construct a minimum spanning tree (MST, Fig. 2(a)) using the Prim algorithm and find a walk that goes around the MST. The solution routing of the Christofides algorithm ensures no longer than twice the optimal routing. However, the Christofides algorithm is designed for symmetric TSP. For exploration-oriented ATSP, the robot does not have to return to the original node as stated by [16]. Therefore, to minimize the routing length, the leaf node that is the farthest from the robot and nodes on its corresponding branch (in Fig. 2(b), the green nodes) will be executed last to save the highest cost of returning to the original node. Another problem is how to decide the sequence of other branches. Since all the regions will be explored finally, the proposed method chooses the next node closest to the current node heuristically when the current node has multiple siblings as shown in Fig. 2(c). Finally, a routing for ATSP will be found as shown in Fig. 2(d).
For the worst case, the path length $l_{eohc}$ of the routing solved by the EOHC algorithm holds:
$l_{eohc}\leq 2l^{*}+l_{m},$ (1)
where $l^{*}$ is the optimal solution of the corresponding ATSP, and $l_{m}$ is the longest distance among all the distances between the robot and all the explorable history nodes. The computational complexity of the proposed EOHC is $\mathcal{O}(n^{3})$, where $n$ is the total number of nodes.
IV-B Curvature-Penalized Viewpoints Planning
After finding the long-term routing of regions, three viewpoints (an exploring viewpoint, a continuous viewpoint, and a safety viewpoint) in the first region of the routing will be selected for ASEO trajectory planning. The UAV will go to explore the exploring viewpoint (in the explored map) and maintain a speed and an acceleration conducive to explore the continuous viewpoint (in the explored map or unexplored map) if no obstacle is detected, otherwise it can return to the safety viewpoint (in the explored map).
For detailed viewpoint evaluation, existing methods usually use the path distance and turning angle to avoid long traveling distances and sharp turns in a decoupled manner. However, the maximum velocity is limited by the curvature of the trajectory at a point according to the circular motion model. We compute the velocity $v_{s}(\mathbf{p})$ at $\mathbf{p}$ by
$v_{s}(\mathbf{p})=\min\left(\sqrt{\frac{\lambda_{a}\cdot a_{m}}{K_{c}(\textbf{p})}},v_{m}\right),$ (2)
where $v_{m}$ is the maximum linear velocity of the UAV, $a_{m}$ is the maximum linear acceleration of the UAV, $K_{c}(\mathbf{p})$ is the curvature of the trajectory at $\mathbf{p}$, and $\lambda_{a}\in(0.0,1.0)$ is an adjustable parameter. Based on this fact, we designed a curvature-penalized heuristic motion cost evaluation method to select the best exploring target and the best continuous target for high-speed exploration.
IV-B1 Curvature-Penalized Exploring Viewpoint and Continuous Viewpoint Planning
An exploring viewpoint in the first region of routing and a continuous viewpoint located in the corresponding EROI of the exploring viewpoint will be selected for aggressive high-speed trajectory planning.
First, a Dijkstra search starts from the current UAV position in the explored space is conducted to find the path to each candidate viewpoint in the first region of the routing. Then, for each path to a candidate viewpoint, an anchor point on the path which is $\lambda_{arc}\cdot v_{m}$ ($\lambda_{arc}$ is a tunable parameter) away from the UAV position is selected to formulate a virtual circular arc trajectory. The formulation of the $i$-th virtual circular arc trajectory to the $k$-th candidate viewpoint $\textbf{vp}_{k}=\left\{\mathbf{p}_{k},\psi_{k}\right\}\in\mathbb{R}^{3}\times SO(2)$ is constrained by the current UAV position $\mathbf{p}_{c}\in\mathbb{R}^{3}$, the current UAV speed $\dot{\mathbf{p}}_{c}\in\mathbb{R}^{3}$, and the anchor point $\mathbf{p}_{a,k}\in\mathbb{R}^{3}$. The center point of the circular arc $\mathbf{p}_{ac,k}\in\mathbb{R}^{3}$ and its curvature can be computed by
$\mathbf{p}_{ac,k}=$
[…]
$K_{c,k}=\frac{1}{\|\mathbf{p}_{ac,k}-\mathbf{p}_{c}\|_{2}}.$ (3)
The virtual velocity $v_{as,k}$ to $\textbf{vp}_{k}$ is penalized by the curvature $K_{c,k}$ and is computed by Eq. 2. The motion cost of going to this exploring target is computed by
$\mathrm{CostE}(\mathbf{vp}_{k})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{c})}{v_{as,k}},\frac{\mathrm{YD}(\psi_{k},\psi_{c})}{\dot{\psi}_{m}}\right),$ (4)
where the function $\mathrm{PL}(\mathbf{p}_{c}$, $\mathbf{p}_{k})$ computes the length of the path from $\mathbf{p}_{c}$ to $\mathbf{p}_{k}$, $\mathrm{YD}(\psi_{k},\psi_{c})$ the yaw difference between the current UAV yaw $\psi_{c}$ and $\psi_{k}$, and $\dot{\psi}_{m}$ is the maximum yaw velocity. Since the terminal velocity at $\mathbf{vp}_{k}$ is unknown, we continue to use the $v_{as,k}$ for computing the continuous motion cost from $\mathbf{vp}_{k}$ to the $j$-th viewpoint $\mathbf{vp}_{j}=\left\{\mathbf{p}_{j},\psi_{j}\right\}\in\mathbb{R}^{3}\times SO(2)$ inside the EROI of $\mathbf{vp}_{k}$
$\mathrm{CostC}(\mathbf{vp}_{j})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{j})}{v_{as,k}},\frac{\mathrm{YD}(\psi_{j},\psi_{k})}{\dot{\psi}_{m}}\right).$ (5)
Finnally, the gain of this viewpoint-pair is evaluated by
$\mathrm{G}(\mathbf{vp}_{k},\mathbf{vp}_{j})=\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{k})}\cdot(1+\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{j})}).$ (6)
Here, $\lambda_{e}$ is a tunable parameter. The larger $\lambda_{e}$ is, the more sensitive it is to the motion costs. The viewpoint-pair with the highest gain is selected as the exploring viewpoint $\mathbf{vp}_{e}$ and the continuous viewpoint $\mathbf{vp}_{c}$.
It is worth noting that the voxels on the continuous line between $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ can not be occupied but free and unknown. This ensures the existence of the possibility for continuous exploration. To efficiently check the feasibilities of the viewpoint pairs, we propose a continuous line-checking method that follows the fast collision checking fashion [36] to avoid recalculation of indexes of voxels at overlapping positions of continuous lines. Specifically, our method generates a table of indexes of voxel that needs to be checked in advance and iterates the table when checking an EROI. Once an iterative voxel is found occupied, its corresponding viewpoint pairs are confirmed infeasible and will not be considered.
IV-B2 Safety Viewpoint Planning
After finding $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$, another viewpoint in the first region of routing will be selected. The safety viewpoint $\mathbf{vp}_{s}$ is the backup target if new obstacles are detected on the trajectory between $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$. To find $\mathbf{vp}_{c}$ efficiently, we directly conduct the Dijkstra algorithm starting from $\mathbf{vp}_{e}$, and the viewpoint that closest to $\mathbf{vp}_{e}$ will be selected as $\mathbf{vp}_{c}$.
Figure 3: An instance of ASEO trajectory. The initial position, velocity, and acceleration of $\mathcal{J}_{e}$ (violet) are fixed with the current UAV position, velocity, and acceleration (yellow). The continuous trajectory $\mathcal{J}_{c}$ (red) ends at $\mathbf{vp}_{c}$. The safety trajectory $\mathcal{J}_{s}$ (red) ends at $\mathbf{vp}_{s}$. The optimizable parameters, including intermediate 4-D points, trajectory durations, velocity, and acceleration at $\mathbf{vp}_{e}$, are illustrated in green.
IV-C ASEO Trajectory Optimization
Inspired by [23], we propose to optimize our 4-D aggressive and safe exploration-oriented (ASEO) trajectory which is based on MINCO trajectory [37]. Different from the 3-D MINCO trajectory, the fourth dimension is the yaw angle, and we optimize it with the position trajectory together.
A 4-D ASEO trajectory $\mathcal{J}(t):\left[t_{s},t_{e}\right]\mapsto\mathbb{R}^{3}\times SO(2)$ ($t_{s}$ and $t_{e}$ are the start time and end time of the trajectory) is paramerized by the initial state $\textbf{S}_{i}=\left\{\mathbf{vp}_{i},\dot{\mathbf{vp}}_{i},\ddot{\mathbf{vp}}_{i}\right\}$, the terminal state $\mathbf{S}_{t}=\left\{\mathbf{vp}_{t},\dot{\mathbf{vp}}_{t},\ddot{\mathbf{vp}}_{t}\right\}$, the $m-1$ intermediate viewpoints $\mathbf{VP}=\left\{\mathbf{vp}_{1},\mathbf{vp}_{2},\cdots,\mathbf{vp}_{m-1}\right\}$, and $m$ durations of trajectory segments $\Delta\mathbf{T}=\left\{\Delta t_{1},\Delta t_{2},\cdots,\Delta t_{m}\right\}$. As shown by Fig. 3, the proposed ASEO trajectory is composed of three connected MINCO trajectories: an exploring trajectory $\mathcal{J}_{e}$ (paramerized by $\textbf{S}_{i,e}$, $\textbf{S}_{t,e}$, $\mathbf{VP}_{e}$, and $\Delta\mathbf{T}_{e}$), a continuous trajectory $\mathcal{J}_{c}$ (paramerized by $\textbf{S}_{i,c}$, $\textbf{S}_{t,c}$, $\mathbf{VP}_{c}$, and $\Delta\mathbf{T}_{c}$), and a safety trajectory $\mathcal{J}_{s}$ (paramerized by $\textbf{S}_{i,s}$, $\textbf{S}_{t,s}$, $\mathbf{VP}_{s}$, and $\Delta\mathbf{T}_{s}$). Some parameters of the ASEO trajectory are fixed. The exploring trajectory starts from the current UAV position $\{\mathbf{p}_{c},\psi_{c}\}$, velocity $\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\}$, and acceleration $\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\}$. The three trajectories are connected at the exploring viewpoint $\mathbf{vp}_{e}$. The continuous trajectory and the safety trajectory end at $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$ correspondingly, their velocity and acceleration are set by zero. We summarize these fixed parameters as:
$\mathbf{vp}_{i,e},\dot{\mathbf{vp}}_{i,e},\ddot{\mathbf{vp}}_{i,e}=\{\mathbf{p}_{c},\psi_{c}\},\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\},\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\},$
$\mathbf{vp}_{t,e}=\mathbf{vp}_{i,c}=\mathbf{vp}_{i,s}=\mathbf{vp}_{e},$
$\mathbf{vp}_{t,c}=\mathbf{vp}_{c},\mathbf{vp}_{t,s}=\mathbf{vp}_{s},$
$\dot{\mathbf{vp}}_{t,c}=\ddot{\mathbf{vp}}_{t,c}=\dot{\mathbf{vp}}_{t,s}=\ddot{\mathbf{vp}}_{t,s}=\{\mathbf{0},0\}.$ (7)
The rest of the parameters are optimizable, and these parameters are shown in Fig. 3 (green). To ensure smoothness, three trajectories share the same velocity and acceleration at the connected point.
The objective function of the ASEO trajectory optimization problem is formulated as follows:
$\textrm{J}_{ASEO}(\mathcal{J}_{e},\mathcal{J}_{c},\mathcal{J}_{s})=j_{e,f}+j_{e,t}+j_{c,f}+j_{c,t}+j_{s,f},$ (8)
where $j_{e,f},j_{c,f}$, and $j_{s,f}$ are the feasibility costs of $\mathcal{J}_{e},\mathcal{J}_{c}$, and $\mathcal{J}_{s}$ corresponingly. To speed up the continuous exploration, $j_{e,t}$ and $j_{c,t}$ (the time costs of $\mathcal{J}_{e}$ and $\mathcal{J}_{c}$) are optimized. The feasibility cost $j_{\mu,f},\mu\in\{e,c,s\}$, it is computed by
$j_{\mu,f}=\int_{t_{s},\mu}^{t_{e,\mu}}\lambda_{p}\mathrm{Vp}_{\mu}(t))+\lambda_{d}(\mathrm{Vv}_{\mu}(t)+\mathrm{Va}_{\mu}(t))dt,$ (9)
where $\mathrm{Vp}_{\mu}(t)$ calculates the position violation of collision-free corridor constraint, $\mathrm{Vv}_{\mu}(t)$ calculates the linear velocity violation of maximum linear velocity constraint, and $\mathrm{Va}_{\mu}(t)$ calculates the maximum acceleration violation of linear acceleration constraint. The penalty weights are $\lambda_{p}$ and $\lambda_{d}$. For the time cost $j_{\mu,t},\mu\in\{e,c\}$, it is computed by
$j_{\mu,f}=\lambda_{t}\sum\limits_{i=1}^{m_{\mu}}\Delta t_{\mu,i},$ (10)
where $t_{\mu,i}\in\Delta T_{\mu}$, $m_{\mu}$ is the number of trajectory segments, and $\lambda_{t}$ is the penalty weight of time. For details of derivation and optimization, we refer the readers to [23].
After generating the ASEO trajectory, $\mathcal{J}_{e}$ is executed. Once $\mathbf{vp}_{e}$ is explored or obstacles are detected on $\mathcal{J}_{c}$, the total planner will plan again.
V experiments and results
V-A Implementation Details
To evaluate the proposed method, we compare our method against state-of-the-art methods [20, 16] in simulation experiments, and conduct real-world experiments.
For comparative simulations, all the methods are implemented in C++ on a computer that runs the Ubuntu 20.04 LTS operation system, it uses Robot Operation System (ROS) for communication between the planner and simulator. The computer is with an Intel Core i7-12700H, a GeForce 3060 12G, and 16G memory. For a fair comparison, all the methods are simulated on the platform that is provided by [20]. The FOV of the RGB-D camera used in the simulation is $[115,92]^{\circ}$, and its maximum sensing range is $5.0\,m$. As for real-world experiments, the proposed method is deployed on a self-made UAV platform (Fig.6 (a)). The proposed method is run on an on-board computer with an Intel Core i7-1260P. The UAV is equipped with a Livox Mid 360 LiDAR for localization and an Intel Realsense D435i RGB-D camera for exploration. The FOV of the Realsense camera is $[87,58]^{\circ}$, and its maximum sensing range is $5.0\,m$. The parameters of our method are shown in Table. I.
TABLE I: Parameters of the proposed method
$r_{v}=0.1m$ $v_{m}=2.0\,m/s$ $a_{m}=3.0\,m/s^{2}$ $\dot{\psi}_{m}=1.57\,rad/s$
$\ddot{\psi}_{m}=1.57\,rad/s^{2}$ $\lambda_{a}=0.6$ $\lambda_{arc}=1.0$ $\lambda_{e}=0.3$
$\lambda_{p}=4000$ $\lambda_{d}=350$ $\lambda_{t}=700$
V-B Simulation Results
In this section, we compare the proposed method against [20, 16] in four simulation environments. Each method runs 10 times in a simulation environment. We record the exploration time, UAV speed, traveled distance, and computational cost for evaluation (shown in Table. II).
  • FALCON [20]: This method plans the exploration sequence of unexplored regions by solving ATSP. After planning the exploration sequence, a time-optimal B-spline trajectory is generated utilizing [38].
  • FUEL [16]: This method plans the exploration sequence of frontiers through solving ATSP. It also uses time-optimal B-spline trajectories for exploration.
TABLE II: EXPLORATION STATISTIC OF ALL METHODS
Environment Method Exploration Time (s) Computational Cost ($ms$) UAV Speed ($m/s$) Traveled Distance ($m$)
Avg Std Avg Std Avg Std Avg Std
DARPA Tunnel Proposed 53.48 1.24 10.31 2.25 1.80 0.018 93.38 2.05
FALCON 56.24 4.50 19.80 6.13 1.63 0.039 96.88 5.84
FUEL 65.32 2.18 10.19 4.59 1.40 0.027 91.78 3.83
Classical Office Proposed 91.56 3.60 14.42 2.94 1.79 0.016 160.86 6.64
FALCON 102.68 3.52 20.00 10.11 1.59 0.046 155.25 6.45
FUEL 128.56 5.71 22.76 18.74 1.31 0.045 167.89 10.65
Large Maze Proposed 301.64 15.65 16.16 3.77 1.86 0.015 557.07 32.13
FALCON 343.84 12.45 63.38 33.03 1.61 0.018 554.98 21.64
FUEL 430.08 13.65 97.81 76.42 1.41 0.011 605.55 25.51
Large Tunnel Proposed 2441.32 51.78 20.48 5.70 3.32 0.017 8095.60 185.38
FALCON - - - - - - - -
FUEL - - - - - - - -
(a) Proposed.
(b) FALCON.
(c) FUEL.
(d) Explored Volume.
(e) Proposed.
(f) FALCON.
(g) FUEL.
(h) Explored Volume.
(i) Velocity Color.
(j) Proposed.
(k) FALCON.
(l) FUEL.
(m) Explored Volume.
Figure 4: The mapping results and the executed trajectory of all the methods. The color of the trajectory denotes the velocity of the UAV. Each row illustrates the simulations in one environment. The right column illustrates the explored volume vs. time.
The first simulation environment is the DARPA Tunnel with a scale of $42\times 20\times 2\,m^{3}$. Each method runs $75\,s$. This environment is relatively simple, so the main factor that affects the exploration efficiency is the speed of the UAV. As shown in Fig.4 (a-c), the proposed method maintains a higher speed than other methods ($10.4\%$ faster than FALCON and $28.6\%$ faster than FUEL). As a result, the proposed method is $5.2\%$ more efficient than FALCON and $16.1\%$ more efficient than FUEL.
The second simulation environment is the Classical Office with a scale of $15\times 30\times 2\,m^{3}$. Each method runs $130\,s$. Compared with DARPA Tunnel, Classical Office is more complex. The walls force the UAV to keep turning which slows down the speed of the UAV. Thanks to the curvature-penalized viewpoints planning and ASEO trajectory, the proposed method is able to avoid sharp turns and generate high-speed trajectories as shown in Fig. 4 (e-g). The UAV speed of the proposed method is $12.6\%$ faster than FALCON’s and $36.6\%$ faster than FUEL’s. Since this environment is more open than DARPA Tunnel, the computational costs grow. The computational cost proposed method is $28.9\%$ lower than FALCON’s and $36.6\%$ lower than FUEL’s. As a result, the proposed method is $12.1\%$ more efficient than FALCON and $40.4\%$ more efficient than FUEL.
The third simulation environment is the Maze with a scale of $40\times 40\times 2\,m^{3}$. Each method runs $500\,s$. As shown by Fig. 4 (j-l), the proposed method keeps high speed in this environment. Because this environment is larger than previous ones, the computational costs of FALCON and FUEL are much higher. The computational cost of the proposed method also grows, but only $12.1\%$ compared to simulation in Classical Office. This is because EOCH can solve ATSP efficiently. The path lengths of the proposed method and FALCON remain similar, which are shorter than FUEL’s, though FALCON and FUEL have better solutions for ATSP. This is because their implicit optimal assumptions fail most of the time. Thus, finding a perfect ATSP solution in exploration is usually not cost-effective. The results show the proposed method’s exploration efficiency is $14.0\%$ more efficient than FALCON and $42.6\%$ more efficient than FUEL.
The last simulation environment is the Large Tunnel with a scale of $251.5\times 327.5\times 2\,m^{3}$. Since this environment is very large, we adopt $v_{m}=4.0m/s$ and $a_{m}=6.0m/s^{2}$ to save the simulation time. The purpose of this trial is to test the computational cost of the proposed method in an extremely large environment. The comparative method FALCON and FUEL fail to run in this environment due to their memory usage exceeding the maximum 16G memory of the computer. On the contrary, the proposed method can accomplish the exploration as shown by Fig. 5. The average computational cost of planning is $20.48\,ms$ which is low enough to plan in real time. The average speed is $3.32\,m/s$. The average exploration time is $2441.32\,s$. The result shows the proposed method is able to explore large-scale environments.
(a) Executed Trajectory and Mapping Result.
(b) Exploration Volume.
Figure 5: Simulation result in Large Tunnel.
(a) UAV Platform.
(b) Executed Trajectory and Mapping Result.
(c) Outdoor Environment.
(d) Exploration Volume.
Figure 6: Experiment results of real-world exploration.
V-C Real-World Experiment
To evaluate the proposed method further, we deploy our method on a self-made UAV (Fig. 6(a)) to explore a $42\times 21\times 3\,m^{3}$ outdoor environment (Fig. 6(c)). The exploration results are shown in Fig. 6 (b) and (d). The exploration cost $145.1\,s$ to travel $260.9\,m$, explore $2520\,m^{3}$ of the environment is explored. During the exploration, the average velocity of the UAV is $1.79m/s$, and the average planning time is $18.5\,ms$.
The simulation results show that the proposed method can plan high-speed exploration trajectories in real-time even in large 3-D environments. The real-world experiments show that the proposed method is also effective.
VI conclusion
In this work, we propose an efficient dual-layer exploration planning method (EDEN). By finding long-term routings through the proposed EOHC algorithm, deciding short-term targets through curvature-penalized evaluation, and generating ASEO trajectories, the proposed method is able to explore large environments efficiently with low computational cost and high speed. The simulation indicates the exploration efficiency of the proposed method is $5.2\%-14.0\%$ higher than state-of-the-art methods. The computational cost of the proposed method is low enough to enable exploration in large-scale environments.